# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ Abstract
In this work, we introduce EMMA-500, a large-scale multilingual language model continue-trained on texts across 546 languages designed for enhanced multilingual performance, with a focus on improving language coverage for low-resource languages. To facilitate continual pre-training, we compile the MaLA corpus, a comprehensive multilingual dataset and enrich it with curated datasets across diverse domains. Leveraging this corpus, we conduct extensive continual pre-training of the Llama 2 7B model, resulting in EMMA-500, which demonstrates robust performance across a wide collection of benchmarks, including a comprehensive set of multilingual tasks and PolyWrite, an open-ended generation benchmark developed in this study. Our results highlight the effectiveness of continual pre-training in expanding large language models’ language capacity, particularly for underrepresented languages, demonstrating significant gains in cross-lingual transfer, task generalization, and language adaptability.
@@
robot Model: huggingface.co/collections/MaLA-LM -¿ EMMA-500
@@ 1 Introduction
Multilingual language models (MLMs) are designed to process and generate text in multiple languages. These models have evolved rapidly over the past decade, fueled by advances in deep learning, e.g., Transformer networks (128), pre-training techniques, and the availability of large-scale multilingual corpora such as mC4 (103) and ROOTS (65). The development of models like BERT (31), GPT, and T5 (103) opened the door for multilingual counterparts such as mBERT, XLM-R (26), mGPT (113), and mT5 (133). These models were trained on massive multilingual corpora, allowing text in dozens of languages to be processed with the same set of model weights. By increasing the scale of both the data and the model, multilingual language models have demonstrated impressive performance on a wide array of tasks, such as text classification, machine translation, and question answering across different languages. A key advantage of multilingual models is their ability to leverage cross-lingual transfer, where knowledge learned from high-resource languages (like English or Chinese) can be applied to other languages. However, despite the remarkable progress in multilingual models, many low-resource languages—those with limited available data—remain underserved. While massive and extensive corpora are readily available for high-resource languages like English, French, and Spanish, languages such as Xhosa and Inuktitut often have only scarce or fragmented data. This disparity between languages tends to make training datasets imbalanced, and multilingual models trained on such imbalanced datasets tend to prioritize high-resource languages, leaving low-resource languages underrepresented.
@@
As the field of MLMs evolves, the role of data becomes increasingly critical in enhancing the performance of the models, particularly when it comes to low-resource languages. To address this need for more and better data, we extend existing work, such as MaLA-500, by expanding the corpus for continual pre-training, coupled with large-scale training methods. We emphasize the creation of a massively multilingual corpus that not only increases the quantity of data but also diversifies the types of texts (e.g., code, books, scientific papers, and instructions). This ensures better language adaptation and broader language coverage, thus improving the representation and performance of multilingual language models, especially for underrepresented languages, ultimately creating more inclusive and versatile language models that cater to a broader linguistic diversity. Our contribution is summarized as follows:
@@
  • We perform continual pre-training with the Llama 2 7B model [^1] Choosing Llama 2 allows us to compare our model with many other models derived from it using continual pre-training. We plan to continue training models based on Llama 3/3.1 in the future. on the multilingual corpus with 546 languages, resulting in the new EMMA-500 model. This model is rigorously evaluated on a diverse set of tasks and benchmarks, including PolyWrite, a novel open-ended generation benchmark developed as part of this work.
@@
  • It contains 939 languages, 546 of which have more than 100k tokens and are used for training our EMMA-500 model, and 74 billion (B) whitespace delimited tokens in total.
@@
  • It comes with four publicly available versions: (1) a noisy version after only basic pre-processing like extraction and harmonization; (2) a cleaned version after data cleaning; (3) a deduplicated version after approximate and exact deduplication; (4) a split version with train and valid splits.
@@
In a comparison with decoder-only LLMs (Section 4) including Llama 2-based continual pre-trained models, LLMs that are designed to be multilingual, and the latest advanced LLMs, our model achieves strong results:
@@
  • Our model remarkably improves the performance of commonsense reasoning, machine translation, and open-ended generation over Llama 2-based models and multilingual baselines, and outperforms the latest advanced models in many cases.
@@ 2.1.1 Extraction and Harmonization
As mentioned, there is significant variability in the data quality of the source datasets used for compiling the corpus. Many of the datasets exhibit data quality challenges that would have adverse effects on model training if left unaddressed. These dataset-specific challenges are typically addressed during the pre-processing stage. For example, we identify one issue involving text records consisting solely of date and timestamp information in the dataset for Languages of Russia (28), likely resulting from a web scraper failing to differentiate between these elements and actual text. We address this by implementing a logic in the pre-processing script to detect and exclude such records. This issue is resolved by introducing a rule to identify and discard text containing consecutive repeating words.
@@ 2.1.3 Writing System Recognition
During script identification, we encounter several challenges. One issue is determining an appropriate length for the text chunk used for script recognition. A chunk that is too short could lead to incorrect identification if the text contains quotes or foreign language fragments using a different writing system. Conversely, using a chunk that is too long could result in excessive resource usage, slowing down processing or even causing memory exhaustion. Another consideration is whether to assume that a single file or dataset might contain multiple scripts. Such an assumption would require identifying the script at a more granular level, such as paragraph by paragraph or even sentence by sentence. Alternatively, we could assume that each file or dataset contained only one “main” script. This assumption would allow us to identify the script from a representative sample of the text for the whole file or dataset. We adopt the latter approach, recognizing a single dominant script for each dataset. The output of this process is a label in the format language_Script, e.g., eng_Latn, where “Language” represents the ISO 639-3 language code and ’‘Script” represents the ISO 15924 script code.
2.1.4 List of Data Sources
The MaLA corpus harvests a wide range of datasets in multiple domains. Table 19 in the appendix lists the corpora and collections we used as monolingual data sources. The major ones include AfriBERTa (95), Bloom library (66), CC100 (26; 129) CulturaX (94), CulturaY (122), the Curse of Multilinguality (21), Evenki Life (70), Glot500 (49), GlotSparse (53), monoHPLT of HPLT v1.2 (30) from the HPLT project (12), Indigenous Languages Corpora (34), Indo4B (131), Lacuna Project (81), Languages of Russia (28), MADLAD-400 (60), Makerere Radio Speech Corpus (89), masakhane-ner1.0 (48), MC2 (137), mC4 (103), multilingual-data-peru (45), OSCAR 2301 (96) from the OSCAR project  [^4] https://oscar-project.org/, Tatoeba challenge monolingual collection (124), The Leipzig Corpora (41), Tigrinya Language Modeling (38), Wikipedia 20231101 (130) and Wikisource 20231201 (130). We exclude high-resource languages in CulturaX, HPLT, MADLAD-400, CC100, mC4, and OSCAR 2301. We exclude Gahuza and Pidgin in the AfriBERTa dataset. We filter out texts that mainly contain a date or timestamp in the Languages of Russia dataset. For Glot500-c, we filter out texts, which may come from train or test tests from datasets for machine translation, such as Flores200, Tatoeba, and mtdata. Despite the translation data being split into source and target languages in the Glot500-c corpus, we decide to filter them to avoid potential data leakage, especially since we use Flores200 as the evaluation benchmark.
@@ 2.2 Data Cleaning
Following the pipeline used by BigScience’s pipeline for ROOTS corpus (65), we further adopt some necessary data cleaning to filter out text samples that might have undesirable quality. We first perform document modification for monolingual texts. The first step is whitespace standardization: all types of whitespace in a document are converted into a single, consistent space character. We split documents by newline characters, tabs, and spaces, strip words, and reconstruct the documents to remove very long words. However, these two steps do not apply to languages without whitespace word delimiters like Chinese, Japanese, Korean, Thai, Lao, Burmese, etc. We also remove words containing certain patterns, e.g., “http” and “.com”, which are likely to be links and page source code. We then perform document filtering, including word count filtering, character repetition filtering, word repetition filtering, special characters filtering, stop words filtering, and flag words filtering. We re-identify the languages that are supported by the pre-trained fastText-based language identification model (51; 50). For other languages, we assume the language identification of the original data source and language code conversion are reliable.
@@ 2.4 Key Statistics
This section presents the final MaLA corpus obtained after data sourcing, pre-processing, cleaning, and deduplication. Table 1 first shows some basic data statistics and compares them with other multilingual corpora for pre-training language models or language adaptation. Additional statistics are presented in Appendix B in the appendix. The token counts are based on white-space delimitation, though it might not be accurate for languages like Chinese, Japanese and Korean since the entire clause is counted as one token. Glot500-c (49) has 534 languages in total, in which 454 languages are directly distributed on Huggingface  [^5] https://huggingface.co/datasets/cis-lmu/Glot500. We also omit high-resource languages in the other three datasets, i.e., MADLAD (60), CulturaX (94), and CC100 (129), as our main focus is continual pre-training for language adaptation. The final MaLA corpus consists of 939 languages, 546 of which have more than 100k tokens and are used for training our EMMA-500 model. Counting languages with more than 1 million tokens, the MaLA corpus and Glot500-c have more than 300, while MADLAD and CulturaX have 200 and 100 respectively. Compared with Glot500-c, the MaLA corpus contains documents with significantly higher sequence lengths, with an average token count of 90 versus 19. This higher sequence length is advantageous for continually training LLMs because it provides more context within each training example, allowing the model to better capture long-range dependencies and patterns in the data. As a result, MaLA is more effective for language adaptation.
@@ 3.1 Data Mixing
We additionally enrich the training corpus by sourcing code data from The Stack (57). This is done following existing work that demonstrates the value of code data in improving the reasoning ability of language models (139; 79) while also mitigating any catastrophic forgetting of the base model’s programming knowledge.
We subsample The Stack at an effective rate of 15.2%, prioritizing high-quality source files and data science code  [^9] https://huggingface.co/datasets/AlgorithmicResearchGroup/arxiv_research_code. We retain the 32 most important non-data programming languages by prevalence while also adding in all llvm code following prior work detailing its importance in multi-lingual code generation (118; 99). We also source from data-heavy formats but follow precedent (75) and subsample them more aggressively. For a more detailed read on filtering heuristics, we direct the reader to Appendix A.2.
Data Mix
Our final data mix for continual training is listed in Table 2. The resource categorization refers to Section B.1 in the appendix and inst medium-high+ is a separate category with languages with more than 500 million but less than 1B tokens. For monolingual text, we also have a seperate category for English. In continual learning, where new data is introduced to an existing model, there is a risk of “catastrophic forgetting”, where the model loses knowledge from earlier training stages. Although our work’s primary focus is in a low-resource regime, we enhance the training corpus with a wide range of data types, including books and scientific papers in high-resource languages, code, and instruction data in our data mix. We downsample texts in high-resource languages and upsample text in low-resource languages using different sample rates according to how resourceful the language is. We make our data mix diverse and balanced towards different resource groups of languages in order to retain the prior knowledge of the model while learning new information, especially in medium- and low-resource languages, thus maintaining high performance across both previously seen and new languages. The final data mix has around 136B tokens.
@@ 4.1 Tasks, Benchmarks, and Baselines
We conduct a comprehensive evaluation to validate the usability of our processed data and data mixing for massively multilingual language adaptation. We perform the intrinsic evaluation of the models’ performance on next-word prediction and evaluate the model’s performance on downstream tasks. Table 3 lists the datasets we used as downstream evaluation datasets in this work. For math tasks, we perform direct prompting and Chain-of-Thoughts prompting on the MGSM benchmark and evaluate the performance using both exact and flexible matches. For code generation tasks, we perform test-case-based execution-tested evaluations using the pass@k metric (22). We benchmark for k values of 1,10, and 25 using a generation pool of 50 samples per problem.
Tasks Dataset Metric Samples/Lang N Lang Domain
Intrinsic Evaluation Glot500-c test (49) NLL 1000 534 Misc
PBC (82) NLL 500 370 Bible
Text Classification SIB200 (6) ACC 204 205 Misc
Taxi1500 (78) ACC 111 1507 Bible
Commonsense Reasoning XCOPA (100) ACC 600 11 Misc
XStoryCloze (73) ACC 1870 11 Misc
XWinograd (125) ACC 741.5 6 Misc
Natural Language Inference XNLI (27) ACC 2490 15 Misc
Machine Translation FLORES-200 (29) BLEU, chrF++ 1012 204 Misc
Open-Ended Generation Aya (115) BLEU, Self-BLEU 215 119 Misc
PolyWrite (Ours) Self-BLEU 149 240 Misc
Summarization XL-Sum (46) ROUGE-L, BERTScore 2537 44 News
Math MGSM direct (112) ACC 250 10 Misc
MGSM CoT (112) ACC 250 10 Misc
Machine Comprehension BELEBELE (14) ACC 900 122 Misc
ARC multilingual (63) ACC 1170 31 Misc
Code Generation Multipl-E (20) Pass@$k$ 164 7 Misc
@@
We compare our model with three groups of decoder-only models. They are (1) Llama 2 models (126) and continual pre-trained models based on Llama 2, such as CodeLlama (106), MaLA-500 (72), LLaMAX (76), Tower (9), and YaYi [^11] https://huggingface.co/wenge-research/yayi-7b-llama2; (2) other LLMs and continual pre-trained LLMs designed to be massively multilingual, including BLOOM (109), mGPT (113), XGLM (73), and Occiglot [^12] https://huggingface.co/occiglot/occiglot-7b-eu5; and (3) recent LLMs with superior English capabilities like Llama 3 (32), Llama 3.1131313https://llama.meta.com/docs/model-cards-and-prompt-formats/llama3_1, Qwen 2 (134), and Gemma 2 (120). There are also some other LLMs such as OpenAI’s API models  [^14] https://platform.openai.com/docs/models and xLLMs-100 (64). However, they do not release the model weights or they limit access to them through commercial API, so we did not include them. The MADLAD model (60) that uses the decoder-only T5 architecture is not supported by inference engines such as the HuggingFace transformers (132). We do not compare them in this work.
Evaluation Software
We use the Language Model Evaluation Harness (lm-evaluation-harness) framework (39) for benchmarking test sets that are already ingested in the framework. For other benchmarks, we use in-house developed evaluation scripts and other open-source implementations. In text classification tasks such as SIB-200 and Taxi-1500, the evaluation protocol involves calculating the probability of the next output token for each candidate category. These probabilities are then sorted in descending order, with the category having the highest probability being selected as the model’s prediction. This process is implemented using the Transformers (132) library. In tasks like Machine Translation and Open-Ended Generation, the vLLM library (62) is used to accelerate inference, providing significant speed improvements. Similarly, for code generation tasks, we use a VLLM-enabled evaluation harness package (vllm-code-harness) [^15] https://github.com/iNeil77/vllm-code-harness for the execution-tested evaluations. For massively multilingual benchmarks with more than 100 languages, we categorize languages into five groups as listed in Table 20 in the appendix.

### v2

@@ Abstract
In this work, we introduce EMMA-500, a large-scale multilingual language model continue-trained on texts across 546 languages designed for enhanced multilingual performance, with a focus on improving language coverage for low-resource languages. To facilitate continual pre-training, we compile the MaLA corpus, a comprehensive multilingual dataset and enrich it with curated datasets across diverse domains. Leveraging this corpus, we conduct extensive continual pre-training of the Llama 2 7B model, resulting in EMMA-500, which demonstrates robust performance across a wide collection of benchmarks, including a comprehensive set of multilingual tasks. Our results highlight the effectiveness of continual pre-training in expanding large language models’ language capacity, particularly for underrepresented languages, demonstrating significant gains in cross-lingual transfer, task generalization, and language adaptability. We release the MaLA corpus, EMMA-500 model weights, scripts, and model generations.
@@
robot Model: huggingface.co/collections/MaLA-LM $\rightarrow$ EMMA-500
@@ 1 Introduction
Multilingual large language models (LLMs) are designed to process and generate text in multiple languages. These models have evolved rapidly over the past decade, fueled by advances in deep learning, e.g., Transformer networks (129), pre-training techniques, and the availability of large-scale multilingual corpora such as mC4 (104) and ROOTS (66). The development of models like BERT (32), GPT, and T5 (104) opened the door for multilingual counterparts such as mBERT, XLM-R (27), mGPT (114), and mT5 (135). These models were trained on massive multilingual corpora, allowing text in dozens of languages to be processed with the same set of model weights. By increasing the scale of both the data and the model, multilingual language models have demonstrated impressive performance on a wide array of tasks, such as text classification, machine translation, and question answering across different languages. A key advantage of multilingual models is their ability to leverage cross-lingual transfer, where knowledge learned from high-resource languages (like English or Chinese) can be applied to other languages. However, despite the remarkable progress in multilingual models, many low-resource languages, which have limited available data, remain underserved. While massive and extensive corpora are readily available for high-resource languages like English, French, and Spanish, languages such as Xhosa and Inuktitut often have only scarce or fragmented data. This disparity between languages tends to make training datasets imbalanced, and multilingual models trained on such imbalanced datasets tend to prioritize high-resource languages, leaving low-resource languages underrepresented.
@@
Figure 1: The number of wins, i.e., the number of times EMMA-500 achieves the best or superior performance compared to other models in the same category across various evaluation tasks and benchmarks. We compare our EMMA-500 Llama 2 7B model to decoder-only LLMs of similar parameter size, including (i) 10 Llama 2-based LLMs, (ii) 7 multilingual LLMs and CPT models, and (iii) 8 recent advanced LLMs (see Section 4.1) on tasks and benchmarks in Table 3. If EMMA-500 scores higher than all compared models on a specific benchmark, it is considered a winning case for that particular evaluation. Our EMMA-500 Llama 2 model outperforms most Llama 2-based, multilingual LLMs and CPT models. Remarkably, our model achieves the best performance on Flores200, Glot500-c, and PBC among all the compared baselines.
As the field of multilingual LLMs evolves, the role of data becomes increasingly critical in enhancing the performance of the models, particularly when it comes to low-resource languages. To address this need for more and better data, we extend existing work, such as MaLA-500, by expanding the corpus for continual pre-training, coupled with large-scale training methods. We emphasize the creation of a massively multilingual corpus that not only increases the quantity of data but also diversifies the types of texts (e.g., code, books, scientific papers, and instructions). This ensures better language adaptation and broader language coverage, thus improving the representation and performance of multilingual language models, especially for underrepresented languages, ultimately creating more inclusive and versatile language models that cater to a broader linguistic diversity. Our contribution is summarized as follows:
@@
  • We perform continual pre-training with the Llama 2 7B model [^1] Choosing Llama 2 allows us to compare our model with many other models derived from it using continual pre-training. We plan to continue training models based on Llama 3/3.1 in the future. on the multilingual corpus with 546 languages, resulting in the new EMMA-500 model. This model is rigorously evaluated on a diverse set of tasks and benchmarks.
@@
  • It contains 939 languages, (546 of which are used for training our EMMA-500 model) and 74 billion (B) whitespace delimited tokens in total.
@@
  • It comes with four publicly available versions: (1) a noisy version after only basic pre-processing like extraction and harmonization; (2) a cleaned version after data cleaning; (3) a deduplicated version after approximate and exact deduplication; (4) a version with train and valid splits.
@@
In a comparison with decoder-only LLMs (Section 4.1) including Llama 2-based continual pre-trained models, LLMs that are designed to be multilingual, and the latest advanced LLMs, our model achieves strong results (Figure 1):
@@
  • Our model remarkably improves the performance of commonsense reasoning, and machine translation over Llama 2-based models and multilingual baselines, and outperforms the latest advanced models in many cases.
@@
Main Achievements
The MaLA corpus marks a significant leap forward in multilingual LLMs, distinguished by its unprecedented scale, diversity, and strategic emphasis on low-resource languages. By integrating a rich spectrum of text types—including code, scientific literature, and instructional data—it establishes a comprehensive and robust foundation for multilingual pretraining. These advancements culminate in EMMA-500, a multilingual LLM that demonstrates enhanced performance across 9 tasks and 15 benchmarks. The scope and impact of this work make it a landmark contribution, particularly in expanding NLP capabilities for underrepresented languages.
@@ 2.1.1 Extraction and Harmonization
As mentioned, there is significant variability in the data quality of the source datasets used for compiling the corpus. Many of the datasets exhibit data quality challenges that would have adverse effects on model training if left unaddressed. These dataset-specific challenges are typically addressed during the pre-processing stage. For example, we identify one issue involving text records consisting solely of date and timestamp information in the dataset for Languages of Russia (29), likely resulting from a web scraper failing to differentiate between these elements and actual text. We address this by implementing a logic in the pre-processing script to detect and exclude such records.
@@ 2.1.3 Writing System Recognition
During script identification, we encounter several challenges. One issue is determining an appropriate length for the text chunk used for script recognition. A chunk that is too short could lead to incorrect identification if the text contains quotes or foreign language fragments using a different writing system. Conversely, using a chunk that is too long could result in excessive resource usage, slowing down processing or even causing memory exhaustion. Another consideration is whether to assume that a single file or dataset might contain multiple scripts. Such an assumption would require identifying the script at a more granular level, such as paragraph by paragraph or even sentence by sentence. Alternatively, we could assume that each file or dataset contained only one “main” script. This assumption would allow us to identify the script from a representative sample of the text for the whole file or dataset. We adopt the latter approach, recognizing a single dominant script for each dataset. The output of this process is a label in the format language_Script, e.g., eng_Latn, where “Language” represents the ISO 639-3 language code and “Script” represents the ISO 15924 script code.
@@ 2.2 Data Cleaning
Following the pipeline used by BigScience’s pipeline for ROOTS corpus (66), we further adopt some necessary data cleaning to filter out text samples that might have undesirable quality. We first perform document modification for monolingual texts. The first step is whitespace standardization: all types of whitespace in a document are converted into a single, consistent space character. We split documents by newline characters, tabs, and spaces, strip words, and reconstruct the documents to remove very long words. However, these two steps do not apply to languages without whitespace word delimiters like Chinese, Japanese, Korean, Thai, Lao, Burmese, etc. We also remove words containing certain patterns, e.g., “http” and “.com”, which are likely to be links and page source code. We then perform document filtering, including word count filtering, character repetition filtering, word repetition filtering, special characters filtering, stop words filtering, and flag words filtering. We re-identify the languages that are supported by the pre-trained fastText-based language identification model (52; 51). For other languages, we assume the language identification of the original data source and language code conversion are reliable.
@@ 2.4 Key Statistics
This section presents the final MaLA corpus obtained after data sourcing, pre-processing, cleaning, and deduplication. Table 1 first shows some basic data statistics and compares them with other multilingual corpora for pre-training language models or language adaptation. Additional statistics are presented in Appendix C in the appendix. The token counts are based on white-space delimitation, though it might not be accurate for languages like Chinese, Japanese and Korean since the entire clause is counted as one token. Glot500-c (50) has 534 languages in total, in which 454 languages are directly distributed on Huggingface  [^4] https://huggingface.co/datasets/cis-lmu/Glot500. We also omit high-resource languages in the other three datasets, i.e., MADLAD (61), CulturaX (95), and CC100 (131), as our main focus is continual pre-training for language adaptation. The final MaLA corpus consists of 939 languages, 546 of which have more than 100k tokens and are used for training our EMMA-500 model. Counting languages with more than 1 million tokens, the MaLA corpus and Glot500-c have more than 300, while MADLAD and CulturaX have 200 and 100 respectively. Compared with Glot500-c, the MaLA corpus contains documents with significantly higher sequence lengths, with an average token count of 90 versus 19. This higher sequence length is advantageous for continually training LLMs because it provides more context within each training example, allowing the model to better capture long-range dependencies and patterns in the data. As a result, MaLA is more effective for language adaptation.
@@ 3.1 Data Mixing
We additionally enrich the training corpus by sourcing code data from The Stack (58). This is done following existing work that demonstrates the value of code data in improving the reasoning ability of language models (141; 80) while also mitigating any catastrophic forgetting of the base model’s programming knowledge.
We subsample The Stack at an effective rate of 15.2%, prioritizing high-quality source files and data science code  [^8] https://huggingface.co/datasets/AlgorithmicResearchGroup/arxiv_research_code. We retain the 32 most important non-data programming languages by prevalence while also adding in all llvm code following prior work detailing its importance in multi-lingual code generation (119; 100). We also source from data-heavy formats but follow precedent (76) and subsample them more aggressively. For a more detailed read on filtering heuristics, we direct the reader to Appendix B.2.
@@
Tasks Dataset Metric Samples/Lang N Lang Domain
Intrinsic Evaluation (Section 4.2) Glot500-c test (50) NLL 1000 534 Misc
PBC (83) NLL 500 370 Bible
Text Classification (Section 4.4) SIB200 (7) ACC 204 205 Misc
Taxi1500 (79) ACC 111 1507 Bible
Commonsense Reasoning (Section 4.5) XCOPA (101) ACC 600 11 Misc
XStoryCloze (74) ACC 1870 11 Misc
XWinograd (126) ACC 741.5 6 Misc
Natural Language Inference (Section 4.6) XNLI (28) ACC 2490 15 Misc
Machine Translation (Section 4.3) FLORES-200 (30) BLEU, chrF++ 1012 204 Misc
Summarization (Section 4.7) XL-Sum (47) ROUGE-L, BERTScore 2537 44 News
Math (Section 4.8) MGSM direct (113) ACC 250 10 Misc
MGSM CoT (113) ACC 250 10 Misc
Machine Comprehension (Section 4.9) BELEBELE (16) ACC 900 122 Misc
ARC multilingual (64) ACC 1170 31 Misc
Code Generation (Section 4.10) Multipl-E (21) Pass@$k$ 164 7 Misc
Data Mix
Our final data mix for continual training is listed in Table 2. The resource categorization refers to Section C.1 in the appendix and inst medium-high+ is a separate category with languages with more than 500 million but less than 1B tokens. For monolingual text, we also have a separate category for English. In continual learning, where new data is introduced to an existing model, there is a risk of “catastrophic forgetting”, where the model loses knowledge from earlier training stages. Although our work’s primary focus is in a low-resource regime, we enhance the training corpus with a wide range of data types, including books and scientific papers in high-resource languages, code, and instruction data in our data mix. We downsample texts in high-resource languages and upsample text in low-resource languages using different sample rates according to how resourceful the language is. We make our data mix diverse and balanced towards different resource groups of languages in order to retain the prior knowledge of the model while learning new information, especially in medium- and low-resource languages, thus maintaining high performance across both previously seen and new languages. The final data mix has around 136B tokens.
@@ 4.1 Tasks, Benchmarks, and Baselines
We conduct a comprehensive evaluation to validate the usability of our processed data and data mixing for massively multilingual language adaptation. We perform the intrinsic evaluation of the models’ performance on next-word prediction and evaluate the model’s performance on downstream tasks. Table 3 lists the datasets we used as downstream evaluation datasets in this work. For math tasks, we perform direct prompting and Chain-of-Thoughts prompting on the MGSM benchmark and evaluate the performance using both exact and flexible matches. For code generation tasks, we perform test-case-based execution-tested evaluations using the pass@k metric (23). We benchmark for k values of 1,10, and 25 using a generation pool of 50 samples per problem. For massively multilingual benchmarks with more than 100 languages, we categorize languages into five groups—high, medium-high, medium, medium-low, and low—based on their token counts in the MaLA corpus as listed in Table 17 in Section C.1.
Evaluation Software
We use the Language Model Evaluation Harness (lm-evaluation-harness) framework (40) for benchmarking test sets that are already ingested in the framework. For other benchmarks, we use in-house developed evaluation scripts and other open-source implementations. In text classification tasks such as SIB-200 and Taxi-1500, the evaluation protocol involves calculating the probability of the next output token for each candidate category. These probabilities are then sorted in descending order, with the category having the highest probability being selected as the model’s prediction. This process is implemented using the Transformers (134) library. In tasks like Machine Translation and Open-Ended Generation, the vLLM library (63) is used to accelerate inference, providing significant speed improvements. Similarly, for code generation tasks, we use a VLLM-enabled evaluation harness package (vllm-code-harness) [^10] https://github.com/iNeil77/vllm-code-harness for the execution-tested evaluations.
@@
We compare our model with three groups of decoder-only models. They are (1) Llama 2 models (127) and continual pre-trained models based on Llama 2, such as CodeLlama (107), MaLA-500 (73), LLaMAX (77), Tower (10), and YaYi [^11] https://huggingface.co/wenge-research/yayi-7b-llama2; (2) other LLMs and continual pre-trained LLMs designed to be massively multilingual, including BLOOM (110), mGPT (114), and Occiglot [^12] https://huggingface.co/occiglot/occiglot-7b-eu5; and (3) recent LLMs with superior English capabilities like Llama 3 (33), Llama 3.1131313https://llama.meta.com/docs/model-cards-and-prompt-formats/llama3_1, Qwen 2 (136), Aya 23 (13) and Gemma 2 (121). There are also some other LLMs such as OpenAI’s API models  [^14] https://platform.openai.com/docs/models and xLLMs-100 (65). However, they do not release the model weights or they limit access to them through commercial API, so we did not include them. The MADLAD model (61) that uses the decoder-only T5 architecture is not supported by inference engines such as the HuggingFace transformers (134). We do not compare them in this work.